A survey of India worker was taking the 2000 census in Behrampur distance of Orissa. It was a Wednesday in the second half of the year in the waxing moon phase right aftera major national festival. In one of the villages the worker intervewed a toothless old lady sitting in front of hut. The lady declared: My grand daughter is about as many days as my daughter is weeks, and my grand daughte is as many months as I am years. The three of us together are 140 years. The surveyor quickly took out her calculator and started punching in a few numbers to find the lady's age. Her calculator has a seven segment LCD displaye but she was puzzled to see the number so she did the calculations again, this time assuming 30 days in a month and 365 days in a year. This time the results were correct. Can you tell what did the calculator display showed, if you made the same assumptions?


a)85 years b)83 years c)84 years d)86 years


 


Grade:11

1 Answers

Aman Bansal
12 years ago

dear doppalapudi,

according to the conditions given in the question,

grand daughter's days= daughter's weeks

grand daughter's months = her years,

let x= age (in years) of old lady herself

y = age (in years) of her daughter

z= age (in years) of her grand daughter

therefore , x+y+z=140    (1)

365z=365/7 * y               (2)

12z=x                             (3)

solving the above three equations we get

x=84 years

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

AMAN BANSAL